# Impact of the latest 22Ne+ reaction rates on nucleosynthesis in massive stars and galactic chemical evolution

### v1

@@ 3.2 Productions of Cu, Zn, Ga, and Ge
While the bulk of the s-process products are expected to be built in pre-SN hydrostatic evolution, the shock wave and explosive nucleosynthesis during CCSN explosions can affect the ejected s-process yields (e.g., Rauscher etĀ al., 2002). Therefore, explosive nucleosynthesis should be taken into account to derive the complete stellar yields. In the following part of this sub-section, we will explore the productions of the elements Cu, Zn, Ga, and Ge, which are located just above the iron-group element Ni, and are produced by the s-process in massive stars (e.g., Raiteri etĀ al., 1991b, a; Rauscher etĀ al., 2002; The etĀ al., 2007; Pignatari etĀ al., 2010). For the remainder of this work, we will focus our analysis and discussions on these elements. As mentioned, it is established that most of the solar Cu, Ga, and Ge are produced by the s-process in massive stars, while the origin of Zn remains a puzzle (e.g., Pignatari etĀ al., 2010; Prantzos etĀ al., 2020; Kobayashi etĀ al., 2020). Thus, it is beneficial to discuss the production processes of these elements in this subsection to better understand their productions in the different layers of the CCSN ejectas, which will lead to better understanding of the impact of the 22Ne+ rates discussed in the later sections.
@@ 3.2.1 Copper
In the solar system, Cu is made up of 69.15% of 63Cu and 30.85% of 65Cu. In Figure 1, we show the abundance profiles for 63Cu in the CCSN ejecta of the 15Mā at =0.02 and =0.0001 models calculated by Ritter etĀ al. (2018b). Note that here denotes the abundance in mass fraction of a given isotope for each mass coordinate in the model222In a given differential mass zone , at the mass coordinate , .. Although the two models have the same initial mass, the model at lower shows a more relevant primary333i.e., not directly affected from the initial composition of the star, in contrast to secondary. explosive nucleosynthesis component, within an -rich region in the deepest 0.2-0.3 Mā region of the ejecta, right above the newly formed proto-neutron star (PNS), where the -process takes place (e.g., Woosley & Hoffman, 1992; Pignatari etĀ al., 2016b). Therefore, we will discuss the primary production of Cu (and other elements discussed in this subsection) in CCSN using the =0.0001 model.
In the model at =0.02, 63Cu shows a mild s-process production in the material ejected from the former convective He shell at mass coordinates between 3.5 Mā and 4.5 Mā. 63Cu at 3.1 Māā3.5 Mā is produced by n-process (Meyer etĀ al., 2000; Pignatari etĀ al., 2018, and references therein) mostly as radiogenic444By radiogenic production, we mean that the given stable isotope, in this case 63Cu, is not ejected directly as itself but in the form of radioactive isotopes that will decay later to 63Cu, based on their decay half-life. 63Ni and 63Co, triggered by the 22Ne(,n)25Mg activation during explosive He burning. The bulk of the s-process production is visible at about 1.8 Māā3 Mā as ashes of the pre-explosive C burning. Traces of more extreme explosive components are present in the most internal ejected layers, but are only barely noticeable in this model.
In the model at =0.0001, the 63Cu production discussed above is about three orders of magnitude smaller, due to the smaller of the progenitor and the well-expected reduction of both the neutron source 22Ne (as it is secondary) and the iron seeds (e.g., Baraffe etĀ al., 1992; Raiteri etĀ al., 1992; Pignatari etĀ al., 2008; Roberti etĀ al., 2024a). Instead, a strong primary production is obtained in the -rich freezout material at 1.6 Māā1.8 Mā. 63Cu show abundances up to a few per mill in mass fractions, and is mostly produced from the radiogenic contribution from 63Zn, with much smaller direct production as 63Cu and the radiogenic contribution from 63Ga.
@@ 3.2.2 Zinc
The solar Zn composition is made up of 49.17% of 64Zn, 27.73% of 66Zn, 4.04% of 67Zn, 18.45% of 68Zn and 0.61% of 70Zn. The nucleosynthesis of Zn in massive stars has been studied extensively for both CCSN and the most energetic hypernovae (Kobayashi etĀ al., 2020, and references therein), due to the puzzling high abundance of Zn with respect to Fe observed in the early Galaxy. This is mostly due to the primary production of 64Zn (and partly 66Zn). In our models, an effective s-process production is obtained for all Zn isotopes except for 64Zn. Since the s-process marginally produces Zn, here we focus in detail on 67Zn (Figure 3) and 68Zn (Figure 4). Analogous figures for 64Zn, 66Zn and 70Zn are available in Appendix.
@@ 3.2.3 Gallium
In the solar system, Ga is made of 69Ga (60.108%) and 71Ga (39.892%). The nucleosynthesis of Ga in CCSN shows remarkable similarities to Cu. Due to the difficulty in observing Ga in stars, a good understanding of their nucleosynthesis production could allow us to use Cu as an observational diagnostic for Ga (Pignatari etĀ al., 2010).
@@ 3.2.4 Germanium
The stable Ge isotopes 70Ge, 72Ge, 73Ge, 74Ge and 76Ge account for 20.52%, 27.45%, 7.76%, 36.52% and 7.75% of the solar Ge, respectively. 70Ge is defined as an s-only isotope, meaning that its production is due to the s-process only, as the r-process contribution is shielded by the stable isobar 70Zn (e.g., Kappeler etĀ al., 2011). In Figure 7, left panel, we confirm this also for the n-process, where there is no direct or radiogenic contribution feeding 70Ge abundances. On the other hand, such a definition may not be entirely accurate, since it is possible to have a primary 70Ge production by the -process, mostly directly making the isotope and with some small radiogenic contribution from 70Se and 70As. A small primary component for Ge is indeed observed in metal-poor stars (Cowan etĀ al., 2005; Pignatari etĀ al., 2010).
@@ 4.1 Nucleosynthesis and stellar yield table - testing the impact of the 22Ne+ rates
As mentioned, we performed the post-processing nucleosynthesis simulations for the 20 massive star progenitor models in the stellar set considered, at the given initial mass ( = 12, 15, 20, and 25 ) and ( = 0.0001, 0.001, 0.006, 0.01, and 0.02). For each progenitor, we performed three calculations using the upper bound, recommended, and lower bound for each of the LO12 and OT20 rates. The upper bound of each reaction rate was probed by using the upper limit of their 22Ne(,n)25Mg rate and the lower limit of their 22Ne(,)26Mg rate. The recommended calculation was made using the recommended rates for both reactions. For the lower bound, we instead used the lower limit of their 22Ne(,n)25Mg rate and the upper limit of their 22Ne(,)26Mg rate. A total of 120 (20 stellar models 6 reaction rates) complete nucleosynthesis simulations were thus performed for LO12 and OT20, including the CCSN nucleosynthesis, to compare with the original yields set by Ritter etĀ al. (2018b).
Additional pre-SN nucleosynthesis calculations are performed for six variants of the OT20 recommended rate, and for the more recently evaluated rates by Adsley etĀ al. (2021) and Wiescher etĀ al. (2023) to better interpret the differences observed between the LO12 and OT20 rates. Thus, an additional 160 nucleosynthesis simulations were performed.
Figure 9 shows the post-SN abundance profiles of 65Cu and 70Ge for the 15 at =0.02 model produced in the present work (similar to Figure 2 and 7 left panels). As expected from the discussion in subsection 3.2, large variations caused by the different 22Ne+ rates are observed mainly between 1.8 and 3 (where the -process is active), for both 65Cu and 70Ge. The largest variation is seen between the high bound setup of LO12 rates and the low bound setup of OT20. Such variations are mostly generated in the convective He-burning core in the progenitor star, carried in the following shell C-burning regions and finally ejected by the CCSN explosion. Regarding the explosive processes discussed earlier, the n-process is not largely affected by 22Ne+ uncertainties, since it operates at temperatures around 1 GK (e.g., Pignatari etĀ al., 2018) and variations between different rates are much smaller than those induced by under He-burning conditions (see e.g., 65Cu in Figure 9).
The process is also not affected by the 22Ne+ rates. On the other hand, uncertainties in the explosion model may significantly affect the appearance of this component, and the use of a more sophisticated CCSN explosion simulation is mandatory to study in detail the nucleosynthesis nearby the PNS (see, e.g., Curtis etĀ al., 2019; Boccioli etĀ al., 2023; Boccioli & Roberti, 2024, and references therein). We plan to address this challenge in future work.
@@ 4.2 GCE Simulations
For AGB stars (with initial masses 1-7), we used the original yields from Ritter etĀ al. (2018b), while for Type Ia SN (SNIa), we used the fast deflagration W7 yields from Iwamoto etĀ al. (1999). We did not perform additional simulations on these astrophysical sources because the productions of Cu, Zn, Ga, and Ge are dominated by massive stars. Likewise, we did not consider the contribution from popIII stars or Hypernovae. These sources could potentially contribute to the abundances of these elements at very low (e.g., Kobayashi etĀ al., 2020). For the present work, we will focus on the GCE of stars with [Fe/H]-2. Stellar yields are interpolated below and above the transition mass of 8 for the AGBs and massive stars, respectively (see Ritter etĀ al., 2018a, for detail).
We used a two-zone GCE code, OMEGA+ (CÓté et al., 2018). The first (inner) zone consists of the star-forming region in the galaxy modeled by OMEGA (CÓté et al., 2017) and the second (outer) zone represents the circumgalactic medium (CGM). The mass fraction of a given isotope () in the gas (with its mass denoted by ) within the inner zone at a given time will be obtained within OMEGA+ as follows (CÓté et al., 2016, 2017). The temporal evolution of the gas within the inner zone () is calculated with the following equation:
@@
where and are the gas coming in or out from external intergalactic medium (IGM), respectively. In the present work, and are set to 0. We tested different constant values for them following the prescription in CÓté et al. (2017) (and references therein; ) and within the range listed in Table 1 of CÓté et al. (2018) () and observed negligible impacts on our GCE results.
At each time-step , OMEGA creates a simple stellar population (SSP) using SYGMA (Ritter etĀ al., 2018a) to model the stellar ejecta. The SSP is a group of stars created at the same time with the same initial metallicity (, zero-age main sequence), where the initial mass function (IMF) of Kroupa (2001) defines the mass distribution of the AGB and massive stars. The stellar ejecta, , at a certain time is calculated by taking the sum of the SSPās ejecta from =1 to the total number of SSPs that have formed by time . Thus,
@@
For other variables (, , and ), we followed a prescription provided by CÓté et al. (2019), in which model parameters were tuned to best reproduce observational constraints for the Milky Way disk (Kubryk et al., 2015).
We used a two-infall model of gas inflow by Chiappini etĀ al. (1997), which accounts for two episodes of exponential gas inflow. This model is described with the following equation:
@@
The values of the parameters used above are taken from CÓté et al. (2019) and Womack et al. (2022), and they are listed in Table 1.
@@
In Figure 10, we plotted some important galactic properties: (SFR), (the gas mass of the galaxy), , , the CCSN rate, and SNIa rate as a function of the galactic age. All these diagnostics fall within the observational limits, reproducing the results from CÓté et al. (2018). Note that the two infall-episodes are not as clearly distinguishable in these plots (e.g., ) as those of Chiappini et al. (1997). The bimodality derived from the two episodes is, however, more clearly visible when using different parameter values as shown in CÓté et al. (2019). Furthermore, since we do not include the threshold in stellar formation process as Chiappini et al. (1997) adopted, the two episodes are more smoothly connected to each other, making the bimodality less apparent.
@@
It should also be noted that, as discussed in CĆ“tĆ© etĀ al. (2016) and CĆ“tĆ© etĀ al. (2019), choice of different and has non-negligible impact on GCE. For instance, when using sets of the parameter values to reproduce the high and low limits of the observational constraints (see Fig. 10 and CĆ“tĆ© etĀ al., 2019), variations of dex for Cu and Zn and dex for Ga and Ge are found in the GCE results below [Fe/H]ā0.5, respectively. However, we confirmed that the relative differences observed in GCE using different reaction rates, which are the main interest of the present work, are negligibly influenced by the choice of the parameter sets. No GCE variations are indeed observed near solar .
@@
Since the values of , the mass-loading constant that controls , are not largely varied in the above parameter sets (0.52, 0.50, and 0.45 for the best-fit, low, and high limits, respectively; defined in Table 1 of CĆ“tĆ© etĀ al. (2019)), we further tested different values for to specifically see the role of in the GCE. The resulting differences in the GCE curves are negligibly small except for the final (or [Fe/H]) where the GCE reaches at the end of the simulation (=13 Gyrs). About 0.1 dex higher (lower) [Fe/H] is reached with a lower (higher) value of =0 (1.0) because the enriched heavy elements in the Galaxy are exchanged for less enriched medium in the CGM. Note that we stop the simulations at =13 Gyrs. The same simulation time was assumed e.g., in Kobayashi etĀ al. (2006), and using the more recently reported Milky Way galaxyās age (13.6 Gyrs) only leads to small differences in our GCE results.
@@
As additional benchmarking, we plotted the temporal evolution of [Fe/H] and in Figure 11. In the figure, reaches the solar value (=0.014 from Asplund etĀ al., 2009) by the time the solar system was formed (8.5 Gyrs) and continues toward moderate supersolar enrichment (=0.02). [Fe/H] however is slightly underestimated (ā0.2 dex) at 8.5 Gyrs, possibly because we used -enhanced initial abundances for stellar yield calculations at 0.006. Since the initial abundances of Fe are relatively small for the corresponding (see above), it could lead to a smaller amount of Fe ejecta from massive stars and SNIa (see e.g., Ritter etĀ al., 2018a). We thus conclude that our GCE simulations are reasonably calibrated and uncertainties derived from the calibration process do not impact the conclusions of this work.
Further details of the GCE code and the adopted GCE setup can be found in CÓté et al. (2018).

### v2

@@ 3.2 Productions of Cu, Zn, Ga, and Ge
While the bulk of the s-process products are expected to be built in pre-SN hydrostatic evolution, the shock wave and explosive nucleosynthesis during CCSN explosions can affect the ejected s-process yields (e.g., rauscher:02). Therefore, explosive nucleosynthesis should be taken into account to derive the complete stellar yields. In the following part of this sub-section, we will explore the productions of the elements Cu, Zn, Ga, and Ge, which are located just above the iron-group element Ni, and are produced by the s-process in massive stars (e.g., raiteri:91; raiteri:91a; rauscher:02; the:07; 2010ApJ...710.1557P). For the remainder of this work, we will focus our analysis and discussions on these elements. As mentioned, it is established that most of the solar Cu, Ga, and Ge are produced by the s-process in massive stars, while the origin of Zn remains a puzzle (e.g., 2010ApJ...710.1557P; Prantzos2020; kobayashi:20). Thus, it is beneficial to discuss the production processes of these elements in this subsection to better understand their productions in the different layers of the CCSN ejectas, which will lead to better understanding of the impact of the 22Ne+ rates discussed in the later sections.
@@ 3.2.1 Copper
In the solar system, Cu is made up of 69.15% of 63Cu and 30.85% of 65Cu. In Figure 1, we show the abundance profiles for 63Cu in the CCSN ejecta of the 15M⊙ at =0.02 and =0.0001 models calculated by Ritter2018. Note that here denotes the abundance in mass fraction of a given isotope for each mass coordinate in the model222In a given differential mass zone , at the mass coordinate , .. Although the two models have the same initial mass, the model at lower shows a more relevant primary333i.e., not directly affected from the initial composition of the star, in contrast to secondary. explosive nucleosynthesis component, within an -rich region in the deepest 0.2-0.3 M⊙ region of the ejecta, right above the newly formed proto-neutron star (PNS), where the -process takes place (e.g., woosley:92; 2016ApJS..225...24P). Therefore, we will discuss the primary production of Cu (and other elements discussed in this subsection) in CCSN using the =0.0001 model.
In the model at =0.02, 63Cu shows a mild s-process production in the material ejected from the former convective He shell at mass coordinates between 3.5 M⊙ and 4.5 M⊙. 63Cu at 3.1 M⊙–3.5 M⊙ is produced by n-process (meyer:00; pignatari:18, and references therein) mostly as radiogenic444By radiogenic production, we mean that the given stable isotope, in this case 63Cu, is not ejected directly as itself but in the form of radioactive isotopes that will decay later to 63Cu, based on their decay half-life. 63Ni and 63Co, triggered by the 22Ne(,n)25Mg activation during explosive He burning. The bulk of the s-process production is visible at about 1.8 M⊙–3 M⊙ as ashes of the pre-explosive C burning. Traces of more extreme explosive components are present in the most internal ejected layers, but are only barely noticeable in this model.
In the model at =0.0001, the 63Cu production discussed above is about three orders of magnitude smaller, due to the smaller of the progenitor and the well-expected reduction of both the neutron source 22Ne (as it is secondary) and the iron seeds (e.g., baraffe:92; raiteri:92; pignatari:08; roberti:24). Instead, a strong primary production is obtained in the -rich freezout material at 1.6 M⊙–1.8 M⊙. 63Cu show abundances up to a few per mill in mass fractions, and is mostly produced from the radiogenic contribution from 63Zn, with much smaller direct production as 63Cu and the radiogenic contribution from 63Ga.
@@ 3.2.2 Zinc
The solar Zn composition is made up of 49.17% of 64Zn, 27.73% of 66Zn, 4.04% of 67Zn, 18.45% of 68Zn and 0.61% of 70Zn. The nucleosynthesis of Zn in massive stars has been studied extensively for both CCSN and the most energetic hypernovae (kobayashi:20, and references therein), due to the puzzling high abundance of Zn with respect to Fe observed in the early Galaxy. This is mostly due to the primary production of 64Zn (and partly 66Zn). In our models, an effective s-process production is obtained for all Zn isotopes except for 64Zn. Since the s-process marginally produces Zn, here we focus in detail on 67Zn (Figure 3) and 68Zn (Figure 4). Analogous figures for 64Zn, 66Zn and 70Zn are available in Appendix.
@@ 3.2.3 Gallium
In the solar system, Ga is made of 69Ga (60.108%) and 71Ga (39.892%). The nucleosynthesis of Ga in CCSN shows remarkable similarities to Cu. Due to the difficulty in observing Ga in stars, a good understanding of their nucleosynthesis production could allow us to use Cu as an observational diagnostic for Ga (2010ApJ...710.1557P).
@@ 3.2.4 Germanium
The stable Ge isotopes 70Ge, 72Ge, 73Ge, 74Ge and 76Ge account for 20.52%, 27.45%, 7.76%, 36.52% and 7.75% of the solar Ge, respectively. 70Ge is defined as an s-only isotope, meaning that its production is due to the s-process only, as the r-process contribution is shielded by the stable isobar 70Zn (e.g., Kappeler2011). In Figure 7, left panel, we confirm this also for the n-process, where there is no direct or radiogenic contribution feeding 70Ge abundances. On the other hand, such a definition may not be entirely accurate, since it is possible to have a primary 70Ge production by the -process, mostly directly making the isotope and with some small radiogenic contribution from 70Se and 70As. A small primary component for Ge is indeed observed in metal-poor stars (cowan:05; 2010ApJ...710.1557P).
@@ 4.1 Nucleosynthesis and stellar yield table - testing the impact of the 22Ne+ rates
As mentioned, we performed the post-processing nucleosynthesis simulations for the 20 massive star progenitor models in the stellar set considered, at the given initial mass ( = 12, 15, 20, and 25 ) and ( = 0.0001, 0.001, 0.006, 0.01, and 0.02). For each progenitor, we performed three calculations using the upper bound, recommended, and lower bound for each of the LO12 and OT20 rates. The upper bound of each reaction rate was probed by using the upper limit of their 22Ne(,n)25Mg rate and the lower limit of their 22Ne(,)26Mg rate. The recommended calculation was made using the recommended rates for both reactions. For the lower bound, we instead used the lower limit of their 22Ne(,n)25Mg rate and the upper limit of their 22Ne(,)26Mg rate. A total of 120 (20 stellar models 6 reaction rates) complete nucleosynthesis simulations were thus performed for LO12 and OT20, including the CCSN nucleosynthesis, to compare with the original yields set by Ritter2018.
Additional pre-SN nucleosynthesis calculations are performed for six variants of the OT20 recommended rate, and for the more recently evaluated rates by Adsley2021 and Wiescher2023 to better interpret the differences observed between the LO12 and OT20 rates. Thus, an additional 160 nucleosynthesis simulations were performed.
Figure 9 shows the post-SN abundance profiles of 65Cu and 70Ge for the 15 at =0.02 model produced in the present work (similar to Figure 2 and 7 left panels). As expected from the discussion in subsection 3.2, large variations caused by the different 22Ne+ rates are observed mainly between 1.8 and 3 (where the -process is active), for both 65Cu and 70Ge. The largest variation is seen between the high bound setup of LO12 rates and the low bound setup of OT20. Such variations are mostly generated in the convective He-burning core in the progenitor star, carried in the following shell C-burning regions and finally ejected by the CCSN explosion. Regarding the explosive processes discussed earlier, the n-process is not largely affected by 22Ne+ uncertainties, since it operates at temperatures around 1 GK (e.g., pignatari:18) and variations between different rates are much smaller than those induced by under He-burning conditions (see e.g., 65Cu in Figure 9).
The process is also not affected by the 22Ne+ rates. On the other hand, uncertainties in the explosion model may significantly affect the appearance of this component, and the use of a more sophisticated CCSN explosion simulation is mandatory to study in detail the nucleosynthesis nearby the PNS (see, e.g., curtis:19; boccioli:23; BR24, and references therein). We plan to address this challenge in future work.
@@ 4.2 GCE Simulations
For AGB stars (with initial masses 1-7), we used the original yields from Ritter2018, while for Type Ia SN (SNIa), we used the fast deflagration W7 yields from Iwamoto1999. We did not perform additional simulations on these astrophysical sources because the productions of Cu, Zn, Ga, and Ge are dominated by massive stars. Likewise, we did not consider the contribution from popIII stars or Hypernovae. These sources could potentially contribute to the abundances of these elements at very low (e.g., kobayashi:20). For the present work, we will focus on the GCE of stars with [Fe/H]-2. Stellar yields are interpolated below and above the transition mass of 8 for the AGBs and massive stars, respectively (see Ritter2018syg, for detail).
We used a two-zone GCE code, OMEGA+ (2018APJ_Cote). The first (inner) zone consists of the star-forming region in the galaxy modeled by OMEGA (2017APJ_Cote) and the second (outer) zone represents the circumgalactic medium (CGM). The mass fraction of a given isotope () in the gas (with its mass denoted by ) within the inner zone at a given time will be obtained within OMEGA+ as follows (Cote2016; 2017APJ_Cote). The temporal evolution of the gas within the inner zone () is calculated with the following equation:
@@
where and are the gas coming in or out from external intergalactic medium (IGM), respectively. In the present work, and are set to 0. We tested different constant values for them following the prescription in 2017APJ_Cote (and references therein; ) and within the range listed in Table 1 of 2018APJ_Cote () and observed negligible impacts on our GCE results.
At each time-step , OMEGA creates a simple stellar population (SSP) using SYGMA (Ritter2018syg) to model the stellar ejecta. The SSP is a group of stars created at the same time with the same initial metallicity (, zero-age main sequence), where the initial mass function (IMF) of 10.1046/j.1365-8711.2001.04022.x defines the mass distribution of the AGB and massive stars. The stellar ejecta, , at a certain time is calculated by taking the sum of the SSP’s ejecta from =1 to the total number of SSPs that have formed by time . Thus,
@@
For other variables (, , and ), we followed a prescription provided by 2019APJ_Cote, in which model parameters were tuned to best reproduce observational constraints for the Milky Way disk (kubryk:15).
We used a two-infall model of gas inflow by 1997ApJ...477..765C, which accounts for two episodes of exponential gas inflow. This model is described with the following equation:
@@
The values of the parameters used above are taken from 2019APJ_Cote and 10.1093/mnras/stac3180, and they are listed in Table 1.
@@
In Figure 10, we plotted some important galactic properties: (SFR), (the gas mass of the galaxy), , , the CCSN rate, and SNIa rate as a function of the galactic age. All these diagnostics fall within the observational limits, reproducing the results from 2018APJ_Cote. Note that the two infall-episodes are not as clearly distinguishable in these plots (e.g., ) as those of 1997ApJ...477..765C. The bimodality derived from the two episodes is, however, more clearly visible when using different parameter values as shown in 2019APJ_Cote. Furthermore, since we do not include the threshold in stellar formation process as 1997ApJ...477..765C adopted, the two episodes are more smoothly connected to each other, making the bimodality less apparent.
@@
It should also be noted that, as discussed in Cote2016 and 2019APJ_Cote, choice of different and has non-negligible impact on GCE. For instance, when using sets of the parameter values to reproduce the high and low limits of the observational constraints (see Fig. 10 and 2019APJ_Cote), variations of dex for Cu and Zn and dex for Ga and Ge are found in the GCE results below [Fe/H]–0.5, respectively. However, we confirmed that the relative differences observed in GCE using different reaction rates, which are the main interest of the present work, are negligibly influenced by the choice of the parameter sets. No GCE variations are indeed observed near solar .
@@
Since the values of , the mass-loading constant that controls , are not largely varied in the above parameter sets (0.52, 0.50, and 0.45 for the best-fit, low, and high limits, respectively; defined in Table 1 of 2019APJ_Cote), we further tested different values for to specifically see the role of in the GCE. The resulting differences in the GCE curves are negligibly small except for the final (or [Fe/H]) where the GCE reaches at the end of the simulation (=13 Gyrs). About 0.1 dex higher (lower) [Fe/H] is reached with a lower (higher) value of =0 (1.0) because the enriched heavy elements in the Galaxy are exchanged for less enriched medium in the CGM. Note that we stop the simulations at =13 Gyrs. The same simulation time was assumed e.g., in Kobayashi2006, and using the more recently reported Milky Way galaxy’s age (13.6 Gyrs) only leads to small differences in our GCE results.
@@
As additional benchmarking, we plotted the temporal evolution of [Fe/H] and in Figure 11. In the figure, reaches the solar value (=0.014 from Asplund2009) by the time the solar system was formed (8.5 Gyrs) and continues toward moderate supersolar enrichment (=0.02). [Fe/H] however is slightly underestimated (–0.2 dex) at 8.5 Gyrs, possibly because we used -enhanced initial abundances for stellar yield calculations at 0.006. Since the initial abundances of Fe are relatively small for the corresponding (see above), it could lead to a smaller amount of Fe ejecta from massive stars and SNIa (see e.g., Ritter2018syg). We thus conclude that our GCE simulations are reasonably calibrated and uncertainties derived from the calibration process do not impact the conclusions of this work.
Further details of the GCE code and the adopted GCE setup can be found in 2018APJ_Cote.